# Mind the Gap: From Resolving Theoretical Foundations of Chiral(ity)-Induced Spin Selectivity to Pioneering Implementations in Quantum Sensing

### v2

@@ I Introduction: On Chirality and Chiral(ity)-Induced Spin Selectivity
The notion of chirality stems from fundamental concepts in geometry. Roughly, an object is chiral if it breaks mirror symmetries along any plane of reflection. In high-energy physics, the Wu experiment famously demonstrated $\beta$-decay to be spin-selective and chiral (in Poincare transformations), establishing how the weak interaction violates parity [1]. Chirality has since come to play a pivotal role across scientific disciplines, for instance in biochemistry—where enantiomers (i.e. chiral images) of the same compounds, that otherwise ought to be chemically identical, can exhibit significant pharmacological differences due to differing chirality [2, 3]. By definition, chiral biomolecules can exist in all their enantiomeric configurations; yet in-vivo, all genetic polymers (DNA/RNA, and their composite residues) exclusively exist in their right-handed enantiomeric forms, while all amino acids are found in the left-handed form. This gives rise to the emergence of self-assembling chiral structures on a macromolecular and cellular scale [4]. Thus, chirality arguably arises as a defining feature of life itself, with living systems by and large exhibiting almost perfect homochirality despite the significant entropic costs entailed [5, 6, 7] (for interesting exceptions, see a review on chirality and extraterrestrial life by Glavin et al. [8]). While such asymmetry strongly suggests that chirality plays a pivotal role in regulating biochemical processes, the precise mechanisms by which it does so remains a long-standing scientific mystery [7, 9]. A prominent candidate to address this gap in our understanding is the phenomenon of chiral(ity)-induced spin-selectivity (CISS), which broadly encompasses how electrons passing through a chiral (and often helical) medium acquire spin-polarization, and other anomalous spin dynamics [10, 11, 12, 13, 14, 15, 16].
@@
A SOC mediated mechanism alone, however, cannot account for the spin-selectivity magnitudes observed in experiments [57, 58, 59, 60], which becomes apparent from an elementary order-of-magnitude [26] estimation. Consider the rest-frame of an electron in motion along a helical path with some incident electric field, $\mathbf{E}_{\mathrm{ext}}$: the electron incurs an effective magnetic field due to the field, $\mathbf{B}_{\mathrm{eff}}=-c^{-2}\vec{v}\times\mathbf{E}_{\mathrm{ext}}$, while possessing a magnetic moment, $\vec{\mu}=-g_{S}\mu_{B}\hbar^{-1}\vec{\sigma}$. The effective magnetic field exerts a torque on the magnetic moment, resulting in a Larmor interaction energy that is equivalent to a Rashba-like spin-orbit coupling: $\Delta H_{\mathrm{SO}}=-\vec{\mu}\cdot\mathbf{B}_{\mathrm{eff}}=\lambda\hat{n}\cdot\left(\vec{\sigma}\times\vec{p}\right)$. By this account, spin-degeneracy is broken and an energy gap opens. Assuming Boltzmann statistics, the energy gap for this spin-splitting gives rise to a population ratio to the factor of $\exp\left(-2\Delta H_{\mathrm{SO}}/{k_{B}T}\right)$. Supposing one expects to retrieve a 20% population ratio at ambient temperatures ($\sim$300 K), this requires an expected spin-splitting field strength of $\sim$20 meV, which is a magnitude apart from expected SOC fields of bare hydrocarbons. Much activity in CISS theory has since been devoted to accounting for the anomalous spin-polarization strengths observed. Concurrently, models attributing CISS’s strength to rich nonequilibrium interfacial effects have also been proposed, from interface-orbital magnetization [61, 62] to spin-blockade & charge-displacement mechanisms [57, 15], collectively giving rise to a more comprehensive theoretical treatment of CISS in nanojunctions.
@@
It is increasingly evident that CISS should not be treated as a single, homogeneous phenomenon, but rather as a set of related effects that manifest differently depending on the experimental context. Each regime—be it solid-state transport, soft-matter photophysics, or biochemical electron transfer—possesses its own assumptions and constraints, requiring tailored theoretical approaches. However, this fragmentation has in turn given rise to a vast array of frameworks that preclude inter-applicability. Given the diversity of materials and platforms in which CISS finds itself relevant, a fundamental question is raised: What lies at the core of CISS? It is our view that, contrary to the perceptible fragmentation in the theoretical literature, a coherent and unifying understanding can nevertheless be gleaned. In this review we proceed by first laying out brief summaries of core mechanisms proposed to underlie CISS. Next, we consolidate developments in the theoretical literature in a systematic manner to give a brief overview of the trends thus far and highlight future directions. We conclude with a discussion of CISS’s technological potential in the context of quantum sensing. We note that this review is not intended as a comprehensive survey of the CISS literature—broader accounts can be found elsewhere [63]—but instead we focus on discerning common identifiable conceptual threads across the diverse manifestations of CISS, the emergence of distinct mechanisms in specific experimental realizations, and the implications these have for emerging quantum metrology applications [64, 65, 22]. We hope this contribution will assist in demystifying CISS and underscore its potential to be primed for near-term applications in spintronics, biochemical industrial application and the quantum information sciences [66].
@@ II.1 Does Spin-Orbit Coupling Alone Give Rise To Spin-Selectivity?
Insofar as SOC is pivotal to CISS, the question of how SOC affects spin-selection is seldom addressed as compared to how SOC may have been enhanced in CISS to close the strength-gap. The spin-momentum locking picture posits simply that electrons with unfavorable spin-states relative to their motion will back-scatter, but by what means will this back-scattering occur? Indeed, Bardarson’s theorem, which follows on from Kramer’s degeneracy, implies that for time-reversal symmetric non-dissipative spin-1/2 systems, the transmission eigenvalues of both spin-configurations should be equal (i.e. spin-polarization should be forbidden in two-terminal junctions) [71]; its implications for CISS have been a chief subject of exploration by Utsumi, Aharony, & Entin-Wohlman [12, 72, 73].
@@
$\delta s=\sum_{k}\langle S\rangle_{k}\gamma_{k}(v_{k}\cdot E)\frac{\partial f_{k}}{\partial E},$ (5)
where $\gamma$ indicates the gamma distribution, and $v$ the spin-wave group velocity. The electric bias also generates a charge current given by:
$j=\frac{e}{V}\sum_{k}v_{k}\delta f_{k},$ (6)
and as such, so long as the spin expectation-value is non-zero for some $k$, it would seem that a charge current is naturally accompanied by a finite spin current. However, is such spin current tantamount to spin-transport or spin-accumulation? Even in the transport regime, it can be shown that such a shifting of the Fermi sea is not equivalent to preferential transmission of spin-species.
We can call upon a physical argument by taking a gauge-covariant formulation of the Hamiltonian with SOC and obtaining its effective electromagnetic fields, an approach initially offered by Fröhlich & Studer [81] and Tokatly [82]. Consider first the generic Pauli-Schrödinger Hamiltonian for a lone electron in an external scalar potential $\pmb{\Phi}$ and vector potential $\mathcal{A}$, with corrections up to $c^{2}$ (derivable from performing the Foldy-Wouthuysen transformation onto the Dirac Hamiltonian):
$\begin{split}H&=\frac{\pmb{\Pi}^{2}}{2m_{0}}+m_{0}c^{2}-e\pmb{\Phi}+\frac{e\hbar}{2m_{0}c}B\cdot\sigma+\frac{e\hbar}{8m_{0}^{2}c^{2}}\mathrm{div}\>E\\ &\quad+\frac{e\hbar}{8m_{0}^{2}c^{2}}[\pmb{\Pi}\cdot(\sigma\times E)+(\sigma\times E)\cdot\pmb{\Pi}],\end{split}$ (7)
@@
$S=\int\mathrm{d}t\,\mathrm{d}\mathbf{r}\>\>i\psi^{\dagger}(D_{0}\psi)-\frac{\hbar}{2m}\sum_{i}(D_{i}\psi)^{\dagger}(D_{i}\psi),$ (12)
@@
$\xi^{\alpha}_{i}$ $=-\frac{i}{2m}\left[\psi^{\dagger}\tau^{\alpha}(D_{i}\psi)-(D_{i}\psi)^{\dagger}\tau^{\alpha}\psi\right]-\frac{1}{4m}A^{\alpha}_{i}\hat{n}$, (14)
@@
In other words, spin-up and spin-down states experience the same dissipative field strength up to a phase difference.
On the contrary, the non-Abelian effective magnetic fields (such as $F_{xy}=\frac{e\hbar}{m_{0}c^{2}}(|E^{x}|-|E^{y}|)\tau^{z}$) may display unequal field strengths between the spinor states. This generates a spin current that is non-dissipative and hence does not contribute in scattering-matrix analyses. Unlike charge currents, non-dissipative spin current can still persist due to the chiral anomaly even if no dissipative spin current can be drawn, and this can result in a magnetic order or spin-texture when coupled with spin-momentum locking, such that biomolecular CISS under bias effectively gives rise to soft matter ferromagnets. Such magnetisation is linked to spontaneous time-reversal symmetry breaking, and is not to be identified with explicit breaking of time-reversal symmetry at the level of the Hamiltonian, which remains invariant [84]. It is thus difficult to take spin-selectivity as an intrinsic consequence of helical geometry. If the spin-filter picture is to be applicable, CISS should rather arise from scattering and external dissipative factors [85, 86], interfacial effects, and charge-trapping [87] etc. as will be discussed in Sec. III.5.
@@ II.2 Onsager’s Relations: Linear or Non-linear? Molecular vs Interfacial Mechanisms?
A persistent question in the CISS corpus is whether anomalous spin-selectivity strengths as reported in experiments originate intrinsically from SOC or whether these are enhanced by medium-specific processes. In other words, can CISS exist as a linear-response phenomenon? Briefly, given an out-of-equilibrium system in the linear regime, for which the coupled flux-force transport equations are $J_{j}=\sum_{k}L_{jk}\mathcal{F}_{k}$ with fluxes $J_{j}$ and generalised forces $F_{k}$, Onsager’s reciprocal relations state that the kinetic coefficients $L_{jk}$ obey $L_{jk}=L_{kj}$ (i.e. they form a symmetric matrix). These relations reflect how microscopic time-reversal symmetry affects macroscopic dynamics, and generally no longer holds when time-reversal symmetry is broken. Accounting for applied fields $B_{m}$ that break time-reversal leads to the generalised Onsager-Casimir relations, $L_{jk}(B_{m})=L_{kj}(B_{m})$.
@@
In the context of two-terminal CISS spin-valve-like set-ups, this means that the charge-conductance measured remains invariant under a reversal of magnetisation (for all magnetised components) and applied magnetic field if present: $G_{ij}(H,M)=G_{ij}(-H,-M)$. The direct implication is that there should be no anisotropy readings for CISS in the linear regime. Since CISS spin-valve experiments clearly do feature magnetocurrent anisotropy, one may conclude that biomolecular-CISS transport observations are inherently nonlinear whenever such setups do not feature the application of an external magnetic field onto target CISS molecules. This also raises the pertinent question: if SOC is not giving rise to both spin-polarization or conductance anisotropy, then what exactly is being captured in readings?
@@
Xiao et al. [87] elaborate on above by comparing CISS to electrical magnetochiral anisotropy (EMCA), in which an external magnetic field is applied to some chiral conductor; both exhibit non-linear response against a driving electric field and feature induced spin-accumulation in leads like Au. Yet, EMCA preserves Onsager-Casimir relations, leading the authors to propose that biomolecular-CISS’s unique magneto-transport properties may arise from charge-trapping within the dielectric bridge. Recent follow-up work by Zhao et al. [87] develops this idea thoroughly, as outlined in Fig. 7: due to the non-Hermitian skin-effect, the wavefunction of the magnetised interface may be exponentially localised either towards the interface or spread further into the bridge, depending on magnetisation direction. This interfacial wavefunction couples with localised impurity/defect-sites in the bridge, and altering the wavefunction distribution also modifies the energy of localised impurity states (effectively, the barrier potential), thus changing the electronic properties of the set-up and circumventing the Onsager-Casimir relations even at zero bias.
@@
$H\cong\frac{\hbar^{2}k_{z}^{2}}{2m}+\alpha^{2}k_{z}\sigma_{z}\cong\frac{\hbar^{2}}{2m}\left(k_{z}\pm\frac{m\alpha_{2}}{\hbar^{2}}\right)^{2},$ (18)

### v3

@@ I Introduction: On Chirality and Chiral(ity)-Induced Spin Selectivity
The notion of chirality stems from fundamental concepts in geometry. Roughly, an object is chiral if it breaks mirror symmetries along any plane of reflection. In high-energy physics, the Wu experiment famously demonstrated $\beta$-decay to be spin-selective and chiral (in Poincaré transformations), establishing how the weak interaction violates parity [1]. Chirality has since come to play a pivotal role across scientific disciplines, for instance in biochemistry—where enantiomers (i.e. chiral images) of the same compounds, that otherwise ought to be chemically identical, can exhibit significant pharmacological differences due to differing chirality [2, 3]. By definition, chiral biomolecules can exist in all their enantiomeric configurations; yet in-vivo, all genetic polymers (DNA/RNA, and their composite residues) exclusively exist in their right-handed enantiomeric forms, while all amino acids are found in the left-handed form. This gives rise to the emergence of self-assembling chiral structures on a macromolecular and cellular scale [4]. Thus, chirality arguably arises as a defining feature of life itself, with living systems by and large exhibiting almost perfect homochirality despite the significant entropic costs entailed [5, 6, 7] (for interesting exceptions, see a review on chirality and extraterrestrial life by Glavin et al. [8]). While such asymmetry strongly suggests that chirality plays a pivotal role in regulating biochemical processes, the precise mechanisms by which it does so remains a long-standing scientific mystery [7, 9]. A prominent candidate to address this gap in our understanding is the phenomenon of chiral(ity)-induced spin-selectivity (CISS), which broadly encompasses how electrons passing through a chiral (and often helical) medium acquire spin-polarization, and other anomalous spin dynamics [10, 11, 12, 13, 14, 15, 16, 17].
@@
A SOC mediated mechanism alone, however, cannot account for the spin-selectivity magnitudes observed in experiments [58, 59, 60, 61], which becomes apparent from an elementary order-of-magnitude [27] estimation. Consider the rest-frame of an electron in motion along a helical path with some incident electric field, $\mathbf{E}_{\mathrm{ext}}$: the electron incurs an effective magnetic field due to the field, $\mathbf{B}_{\mathrm{eff}}=-c^{-2}\vec{v}\times\mathbf{E}_{\mathrm{ext}}$, while possessing a magnetic moment, $\vec{\mu}=-g_{S}\mu_{B}\hbar^{-1}\mathbf{S}$. The effective magnetic field exerts a torque on the magnetic moment, resulting in a Larmor interaction energy that is equivalent to a Rashba-like spin-orbit coupling: $\Delta H_{\mathrm{SO}}=-\vec{\mu}\cdot\mathbf{B}_{\mathrm{eff}}=\lambda\hat{n}\cdot\left(\vec{\sigma}\times\vec{p}\right)$. By this account, spin-degeneracy is broken and an energy gap opens. Assuming Boltzmann statistics, the energy gap for this spin-splitting gives rise to a population ratio to the factor of $\exp\left(-2\Delta H_{\mathrm{SO}}/{k_{B}T}\right)$. Supposing one expects to retrieve a 20% polarization at ambient temperatures ($\sim$300 K), this requires an expected spin-splitting field strength of $\sim$5 meV, which is magnitudes apart from expected SOC splitting energies of bare hydrocarbons. Much activity in CISS theory has since been devoted to accounting for the anomalous spin-polarization strengths observed. Concurrently, models attributing CISS’s strength to rich nonequilibrium interfacial effects have also been proposed, from interface-orbital magnetization [62, 63] to spin-blockade & charge-displacement mechanisms [58, 16], collectively giving rise to a more comprehensive theoretical treatment of CISS in nanojunctions.
@@
It is increasingly evident that CISS should not be treated as a single, homogeneous phenomenon, but rather as a set of related effects that manifest differently depending on the experimental context. Each regime—be it solid-state transport, soft-matter photophysics, or biochemical electron transfer—possesses its own assumptions and constraints, requiring tailored theoretical approaches. However, this fragmentation has in turn given rise to a vast array of frameworks that preclude inter-applicability. Given the diversity of materials and platforms in which CISS finds itself relevant, a fundamental question is raised: What lies at the core of CISS? It is our view that, contrary to the perceptible fragmentation in the theoretical literature, a coherent and unifying understanding can nevertheless be gleaned. In this review we proceed by first laying out brief summaries of core mechanisms proposed to underlie CISS. Next, we consolidate developments in the theoretical literature in a systematic manner to give a brief overview of the trends thus far and highlight future directions. We conclude with a discussion of CISS’s technological potential in the context of quantum sensing. We note that this review is not intended as a comprehensive survey of the CISS literature—broader accounts can be found elsewhere [64]—but instead we focus on discerning common identifiable conceptual threads across the diverse manifestations of CISS, the emergence of distinct mechanisms in specific experimental realizations, and the implications these have for emerging quantum metrology applications [65, 66, 23]. We hope this contribution will assist in demystifying CISS and underscore its potential to be primed for near-term applications in spintronics, biochemical industrial application and the quantum information sciences [67, 68, 69].
@@ II.1 Does Spin-Orbit Coupling Alone Give Rise To Spin-Selectivity?
Insofar as SOC is pivotal to CISS, the question of how SOC affects spin-selection is seldom addressed as compared to how SOC may have been enhanced in CISS to close the strength-gap. The spin-momentum locking picture posits simply that electrons with unfavorable spin-states relative to their motion will back-scatter, but by what means will this back-scattering occur? Indeed, Bardarson’s theorem, which follows on from Kramers’ degeneracy, implies that for time-reversal symmetric non-dissipative spin-1/2 systems, the transmission eigenvalues of both spin-configurations should be equal (i.e. spin-polarization should be forbidden in two-terminal junctions) [74]; its implications for CISS have been a chief subject of exploration by Utsumi, Aharony, & Entin-Wohlman [12, 75, 76].
@@
$\delta s=\sum_{k}\langle S\rangle_{k}\gamma^{-1}_{k}(v_{k}\cdot E)\frac{\partial f^{0}_{k}}{\partial\varepsilon_{k}},$ (5)
where $\gamma_{k}$ indicates the scattering rate, $\varepsilon_{k}$ the band energy, and $v_{k}$ the spin-wave group velocity of the k-th mode. The electric bias also generates a charge current given by:
$j=-\frac{e}{V}\sum_{k}v_{k}\delta f_{k},$ (6)
and as such, so long as the spin expectation-value is non-zero for some $k$, it would seem that a charge current is naturally accompanied by a finite spin current. However, is such spin current tantamount to spin-transport or spin-accumulation? Even in the transport regime, it can be shown that such a shifting of the Fermi sea is not equivalent to preferential transmission of spin-species. We can call upon a physical argument by taking a gauge-covariant formulation of the Hamiltonian with SOC and obtaining its effective electromagnetic fields, an approach initially offered by Fröhlich & Studer [84] and Tokatly [85]. Consider first the generic Pauli-Schrödinger Hamiltonian for a lone electron in an external scalar potential $\pmb{\Phi}$ and vector potential $\mathcal{A}$, with corrections up to $c^{2}$ (derivable from performing the Foldy-Wouthuysen transformation onto the Dirac Hamiltonian):
$\begin{split}H&=\frac{\pmb{\Pi}^{2}}{2m_{0}}+m_{0}c^{2}-e\pmb{\Phi}+\frac{e\hbar}{2m_{0}c}B\cdot\sigma+\frac{e\hbar^{2}}{8m_{0}^{2}c^{2}}\mathrm{div}\>E\\ &\quad+\frac{e\hbar}{8m_{0}^{2}c^{2}}[\pmb{\Pi}\cdot(\sigma\times E)+(\sigma\times E)\cdot\pmb{\Pi}],\end{split}$ (7)
@@
$S=\int\mathrm{d}t\,\mathrm{d}\mathbf{r}\>\>i\hbar\psi^{\dagger}(D_{0}\psi)-\frac{\hbar^{2}}{2m_{0}}\sum_{i}(D_{i}\psi)^{\dagger}(D_{i}\psi),$ (12)
@@
$\xi^{\alpha}_{i}$ $=-\frac{i}{2m_{0}}\left[\psi^{\dagger}\tau^{\alpha}(D_{i}\psi)-(D_{i}\psi)^{\dagger}\tau^{\alpha}\psi\right]-\frac{1}{4m_{0}}A^{\alpha}_{i}\hat{n}$, (14)
@@
In other words, spin-up and spin-down states experience the same dissipative field strength up to a phase difference. On the contrary, the non-Abelian effective magnetic fields (such as $F_{xy}=\frac{e\hbar}{m_{0}c^{2}}(|E^{x}|-|E^{y}|)\tau^{z}$) may display unequal field strengths between spinor states. This generates a spin current that is non-dissipative and hence does not contribute in scattering-matrix analyses. Unlike charge currents, non-dissipative spin current can still persist due to chiral anomaly even if no dissipative spin current can be drawn, and this can result in a magnetic order or spin-texture when coupled with spin-momentum locking, such that biomolecular CISS under bias effectively gives rise to soft matter ferromagnets. Such magnetisation is linked to spontaneous time-reversal symmetry breaking, and is not to be identified with explicit breaking of time-reversal symmetry at the level of the Hamiltonian, which remains invariant [87]. It is thus difficult to take spin-selectivity as an intrinsic consequence of helical geometry. If the spin-filter picture is to be applicable, CISS should arise from scattering and external dissipative factors [88, 89], interfacial effects, and charge-trapping [90] etc. as will be discussed in Sec. III.5.
@@ II.2 Onsager’s Relations: Linear or Non-linear? Molecular vs Interfacial Mechanisms?
A persistent question in the CISS corpus is whether anomalous spin-selectivity strengths as reported in experiments originate intrinsically from SOC or whether these are enhanced by medium-specific processes. In other words, can CISS exist as a linear-response phenomenon? Briefly, given an out-of-equilibrium system in the linear regime, for which the coupled flux-force transport equations are $J_{j}=\sum_{k}L_{jk}\mathcal{F}_{k}$ with fluxes $J_{j}$ and generalised forces $F_{k}$, Onsager’s reciprocal relations state that the kinetic coefficients $L_{jk}$ obey $L_{jk}=L_{kj}$ (i.e. they form a symmetric matrix). These relations reflect how microscopic time-reversal symmetry affects macroscopic dynamics, and generally no longer holds when time-reversal symmetry is broken. Accounting for applied fields $B_{m}$ that break time-reversal leads to the generalised Onsager-Casimir relations, $L_{jk}(B_{m})=L_{kj}(-B_{m})$.
@@
In the context of two-terminal CISS spin-valve-like set-ups, this means that the charge-conductance measured remains invariant under a reversal of magnetisation (for all magnetised components) and applied magnetic field if present: $G_{ij}(H,M)=G_{ij}(-H,-M)$. The direct implication is that there should be no anisotropy readings for CISS in the linear regime. Since CISS spin-valve experiments clearly do feature magnetocurrent anisotropy, one may conclude that biomolecular-CISS transport observations are inherently nonlinear whenever such setups do not feature the application of an external magnetic field onto target CISS molecules. This also raises the pertinent question: if SOC is not giving rise to both spin-polarization or conductance anisotropy, then what is being captured in readings?
@@
Zhao et al. [90] elaborate on above by comparing CISS to electrical magnetochiral anisotropy (EMCA), in which an external magnetic field is applied to some chiral conductor; both exhibit non-linear response against a driving electric field and feature induced spin-accumulation in leads like Au. Yet, EMCA preserves Onsager-Casimir relations, leading the authors to propose that biomolecular-CISS’s unique magneto-transport properties may arise from charge-trapping within the dielectric bridge. As outlined in Fig. 7, due to the non-Hermitian skin-effect, the wavefunction of the magnetised interface may be exponentially localised either towards the interface or spread further into the bridge, depending on magnetisation direction. This interfacial wavefunction couples with localised impurity/defect-sites in the bridge, and altering the wavefunction distribution also modifies the energy of localised impurity states (effectively, the barrier potential), thus changing the electronic properties of the set-up and circumventing the Onsager-Casimir relations even at zero bias.
@@
$H\cong\frac{\hbar^{2}k_{z}^{2}}{2m}+\alpha k_{z}\sigma_{z}\cong\frac{\hbar^{2}}{2m}\left(k_{z}\pm\frac{m\alpha}{\hbar^{2}}\right)^{2},$ (18)